A metallurgist needs to make 3 oz. of an alloy containing 76% platinum. She is going to melt and combine one metal that is 90% platinum with another metal that is 70% platinum. How much of each should she use?
Answer by mananth(16946)

percent percentage ---------------- quantity
Metal I 90 ---------------- x Oz
Metal II 70 ---------------- 3 - x Oz
Mixture 76.00% ---------------- 3

90x+70( 3-x) = 76 * 3

90x+210 - 70 x = 228
90 x - 70 x= 228 - 210
20 x = 18
/ 20
x = 0.9Oz 90.00% Metal I
2.1 Oz 70.00% Metal II